# When isometry and equivalence for skew constacyclic codes coincide

### v1

@@ 2. Petit algebras and power associativity
$t^{i}t^{j}=\begin{cases}t^{i+j}&\text{if $i+j<\mathbf{n}$;}\\ t^{i+j-\mathbf{n}}f_{0}(t)&\text{if $i+j>\mathbf{n}$.}\end{cases}$
Note that $S[t;\sigma]/S[t;\sigma]f(t)$ is associative exactly when $S[t;\sigma]f(t)$ is a two-sided ideal. For example, this happens when $S=K$, $f(t)=t^{\mathbf{n}}-\sum_{i=0}^{\mathbf{n}-1}a_{i}t^{i}\in S_{0}[t]$ and $\sigma^{\mathbf{n}-i}=\operatorname{id}$ for all those $i$ where $a_{i}\not=0$ [Pet67, (15)].
We are particularly interested in the special case that $f(t)=t^{\mathbf{n}}-a$ for some $a\in S^{\times}$. Then the lack of power associativity is best understood by observing that in $S[t;\sigma]$, we have the identity $t^{\mathbf{n}+1}=t^{\mathbf{n}}t=tt^{\mathbf{n}}$ but in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$,
@@
That is, one must take care to simplify products only with right division by $t^{\mathbf{n}}-a$, as then it is clear that the correct reduction is
$t^{\mathbf{n}+1}\equiv t(t^{\mathbf{n}}-a+a)\equiv ta\equiv\sigma(a)t\text{ in }S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a).$
From this point onwards we write simply equality rather than congruence.
We require some additional notation. For any $i\in\mathbb{N}$, $\tau\in\operatorname{Aut}(S)$ and $\beta\in S$, write
$N_{i}^{\tau}(\beta)=\prod_{j=0}^{i-1}\tau^{j}(\beta).$
Note that $\beta\tau(N_{i}^{\tau}(\beta))=N_{i+1}^{\tau}(\beta)$ and thus
(2.1) $N_{i+j}^{\tau}(\beta)=N_{i}^{\tau}(\beta)\cdot\sigma^{i}(N_{j}^{\tau}(\beta))$
for all $i,j\geq 0$. If $S=K$ is a field, $K/K_{0}$ a cyclic Galois extension, and $\operatorname{Gal}(K/K_{0})=\langle\sigma\rangle$, where $\sigma$ has order $\mathbf{m}$, then $N_{\mathbf{m}}^{\sigma}$ is simply the norm map $N_{K/K_{0}}$ corresponding to $K/K_{0}$. More generally, $N_{\mathbf{m}}^{\sigma}(\beta)\in S_{0}$ for all $\beta\in S$ so we distinguish this case with the notation $N_{\mathbf{m}}^{\sigma}=N_{S/S_{0}}$.
Let us define the class of homomorphisms of Petit algebras that correspond to Hamming-weight preserving homomorphisms of skew polycyclic codes.
@@
Let $f,h\in S[t;\sigma]$ be monic polynomials of degree $\mathbf{n}$. Suppose $\tau\in\operatorname{Aut}(S)$, $\alpha\in S$ is nonzero and $k\in\mathbb{N}$. For $R=S[t;\sigma]$ or $R=S[t;\sigma]/S[t;\sigma]f(t)$, if there exists a ring homomorphism
$G_{\tau,\alpha,k}:R\to S[t;\sigma]/S[t;\sigma]h(t)$
defined via $G_{\tau,\alpha,k}|_{S}=\tau$ and $G_{\tau,\alpha,k}(t)=\alpha t^{k}$, that is,
(2.2) $G_{\tau,\alpha,k}\left(\sum_{i=0}^{\mathbf{n}-1}a_{i}t^{i}\right)=\sum_{i=0}^{\mathbf{n}-1}\tau(a_{i})(\alpha t^{k})^{i}.$
then we call it a monomial homomorphism of degree $k$. When $k=1$, we write $G_{\tau,\alpha}$ in place of $G_{\tau,\alpha,1}$, in which case we have simply
@@
In this paper, we expand the set of cases for which it is known that all homomorphisms or isomorphisms of Petit algebras are monomial. This was previously known in several cases, but so far only for isomorphisms.
Then the following results were originally stated for $S_{0}$-isomorphisms for the case that $S$ is a field in [BP18, Theorems 28, 29], and generalized the present setting in [Pum25]. To state them, set
@@
Let $\alpha\in S^{\times}$ and assume that $\tau\in{\rm Aut}(S)$ commutes with $\sigma$. Then
$G_{\tau,\alpha}:S[t;\sigma]/S[t;\sigma]f(t)\to S[t;\sigma]/S[t;\sigma]h(t)$
is an isomorphism of $S_{0}^{\tau}$-algebras if and only if
@@
When $K[t;\sigma]/K[t;\sigma]f(t)$ and $K[t;\sigma]/K[t;\sigma]h(t)$ are proper nonassociative, and $\operatorname{Aut}(K)$ is abelian, we can say even more.
@@
Let $\mathbf{m}\geq\mathbf{n}-1$ and assume that $f$ and $h$ do not generate two-sided ideals in $K[t;\sigma]$. Suppose $\operatorname{Aut}(K)$ is abelian. For any subfield $F\subset K_{0}$, the only $F$-algebra isomorphisms
$G:K[t;\sigma]/K[t;\sigma]f(t)\to K[t;\sigma]/K[t;\sigma]h(t)$
are of the form $G=G_{\tau,\alpha}$ such that $\tau\in\operatorname{Aut}(K)$, $K_{0}^{\tau}\subset F$, and $\alpha\in K^{\times}$ satisfying Theorem 2.2.
@@
3. Monomial homomorphisms from $S[t;\sigma]$ to a Petit algebra
@@
Let $S$ be a commutative unital ring. We begin with general restriction on monomial homomorphisms $G:S[t;\sigma]\to S[t;\sigma]$.
@@
The cancellability of $\alpha$ implies that this holds for all $c\in S$ if and only if $\tau\sigma=\sigma^{k}\tau$. In particular, if $\sigma$ and $\tau$ commute, we must have $\sigma=\sigma^{k}$, or $k\equiv 1\mod\mathbf{m}$. ∎
A special case of this result has appeared as early as [Rim78, Theorem 3].
@@
Thus $G$ induces an isomorphism of $S_{0}^{\tau}$-algebras
$G:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)\quad\text{where $\tau(a)=bN^{\sigma}_{\mathbf{n}}(\alpha)$.}$
In this section we show that these are all the Hamming-weight preserving homomorphisms from $S[t;\sigma]$ to $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$.
The first key obstruction for a map into $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ to be a homomorphism is the power-associativity of the term $\alpha t^{k}$. We require a technical lemma that will in Theorem 3.3 be used to reduce the question of associativity of powers of $\alpha t^{k}$ in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ to a straightforward relation between $\alpha$ and $b$. Write $\lfloor x\rfloor$ for the greater integer that is less than or equal to $x$.
@@
We now demonstrate that the condition (3.1) is exactly the key to the power associativity of a monomial in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$.
@@
in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$. If $\mathsf{z}^{r}\mathsf{z}=\mathsf{z}\mathsf{z}^{r}$ for the least $r>0$ such that $rk\geq{\mathbf{n}}$, then
@@
Conversely, if (3.3) holds, then for every $s\in\mathbb{N}$, the expression $(\alpha t^{k})^{s}$ is well-defined in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ (that is, $\mathsf{z}$ is power associative) and we have
@@
We now prove that if $\mathsf{z}\mathsf{z}^{r}=\mathsf{z}^{r}\mathsf{z}$, for $r>0$ the least integer such that $rk\geq\mathbf{n}$, then (3.3) holds. We begin by noting that $\mathsf{z}^{r}$ is well-defined, since for all $s,\ell\geq 1$ such that $s+\ell=r$,
@@
In particular, Theorem 3.3 shows that the condition (3.3) is both necessary and sufficient for full power associativity of the element $\alpha t^{k}$ in $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$.
@@
$G=G_{\tau,\alpha,k}:S[t;\sigma]\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$
given by $G(a)=\tau(a)$ and $G(t)=\alpha t^{k}$ induces a homomorphism of $S_{0}^{\tau}$-algebras if and only if
@@
Since $\tau$ commutes with $\sigma$ this holds if and only if $s-[sk]_{\mathbf{n}}\equiv 0\mod\mathbf{m}$; since $k\equiv 1\mod\mathbf{m}$ this implies $\mathbf{m}|\mathbf{n}$. Under this hypothesis, the condition $\alpha b=\sigma^{\mathbf{n}}(\alpha)\sigma^{k}(b)$ simplifies to $\sigma(b)=b$ or $b\in S_{0}^{\times}$, and thus $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is associative. Consequently the function $G=G_{\tau,\alpha,k}$ descends to a well-defined homomorphism $G:S[t;\sigma]\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$. ∎
@@
These conditions are quite restrictive, reflecting the constrained nature of these nonassociative algebras. For example, one deduces that even with $k=1$, the map $G_{\tau,\alpha,1}$ is a homomorphism from $S[t;\sigma]\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ only when $b\in S_{0}^{\times}$ and $\mathbf{m}|\mathbf{n}$. This is in fact to be expected: $S[t;\sigma]$ is associative whereas $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is associative if and only if these two strong conditions hold [Pet67, (15)].

### v3

@@ 2.1. Petit rings are not quotient rings
(2.1) $t^{i}t^{j}=\begin{cases}t^{i+j}&\text{if $i+j<\mathbf{n}$;}\\ t^{i+j-\mathbf{n}}f_{0}(t)&\text{if $i+j\geq\mathbf{n}$.}\end{cases}$
2.2. The element $t$ is not always power associative
We are particularly interested in the special case that $f(t)=t^{\mathbf{n}}-a$ for some $a\in S^{\times}$. One must take care to simplify products only with right division by $t^{\mathbf{n}}-a$. A key observation from (2.1) is that in $\mathbb{S}_{t^{\mathbf{n}}-a}$ we have
@@ 2.2. The element $t$ is not always power associative
This example shows that when $a\not\in S_{0}$, we have
$t^{\mathbf{n}}t\not=tt^{\mathbf{n}}$
even though $t^{i}t^{j}=t^{i+j}$ for $i+j<\mathbf{n}$, by (2.1) (with $f_{0}(t)=a$). In general, the element $t$ is power-associative in the ring $\mathbb{S}_{f}$ (which is equivalent to saying all powers of $t$ commute), if and only if $t^{\mathbf{n}}t=tt^{\mathbf{n}}$, which by (2.1) is equivalent to $f(t)t\in S[t;\sigma]f(t)$; this, in turn, is equivalent to $t$ lying in the right nucleus of $\mathbb{S}_{f}$ [Pet67, (5), p. 13-06].
Our next concern is to consider the class of homomorphisms between Petit rings that will induce Hamming-weight preserving maps between the corresponding $(\mathbf{n},\sigma)$-skew polycyclic codes (Section 2.4).
2.3. Monomial homomorphisms between Petit rings
We require some additional notation. For any $i\in\mathbb{N}$, $\sigma\in\operatorname{Aut}(S)$ and $\beta\in S$, write
$N_{i}^{\sigma}(\beta)=\prod_{j=0}^{i-1}\sigma^{j}(\beta).$
Note that $\beta\sigma(N_{i}^{\sigma}(\beta))=N_{i+1}^{\sigma}(\beta)$ and thus
(2.2) $N_{i+j}^{\sigma}(\beta)=N_{i}^{\sigma}(\beta)\cdot\sigma^{i}(N_{j}^{\sigma}(\beta))$
for all $i,j\geq 0$ (as also observed in [CherLer2016, Proposition 2.1] for the case of finite fields). If $S=K$ is a field, $K/K_{0}$ a cyclic Galois extension, and $\operatorname{Gal}(K/K_{0})=\langle\sigma\rangle$, where $\sigma$ has order $\mathbf{m}$, then $N_{\mathbf{m}}^{\sigma}$ is simply the norm map $N_{K/K_{0}}$ corresponding to $K/K_{0}$. Even when $S$ is not a field, under these hypotheses $N_{\mathbf{m}}^{\sigma}(\beta)\in S_{0}$ for all $\beta\in S$ so we distinguish this case with the notation $N_{\mathbf{m}}^{\sigma}=N_{S/S_{0}}$.
Let us define the class of homomorphisms between Petit rings that we will show in Section 2.4 correspond to Hamming-weight preserving homomorphisms of skew polycyclic codes.
@@ 2.3. Monomial homomorphisms between Petit rings
Let $f,h\in S[t;\sigma]$ be monic polynomials of degree $\mathbf{n}$. Suppose $\tau\in\operatorname{Aut}(S)$, $\alpha\in S$ is nonzero and $k\in\mathbb{N}$. If there exists a ring homomorphism
$G:\mathbb{S}_{f}\to\mathbb{S}_{h}$
that is defined via $G|_{S}=\tau$ and $G(t)=\alpha t^{k}$, that is,
(2.3) $G\left(\sum_{i=0}^{\mathbf{n}-1}a_{i}t^{i}\right)=\sum_{i=0}^{\mathbf{n}-1}\tau(a_{i})(\alpha t^{k})^{i}.$
then we call it a monomial homomorphism of degree $k$ and write $G=G_{\tau,\alpha,k}$. When $k=1$, we write $G_{\tau,\alpha}$ in place of $G_{\tau,\alpha,1}$, in which case we have simply
@@
By abuse of notation, we will later denote ring homomorphism $G:R\to R$ or $G:R\to\mathbb{S}_{h}$ that are also defined via $G|_{S}=\tau$ and $G(t)=\alpha t^{k}$ by $G_{\tau,\alpha,k}$ as well.
@@
In this paper, we expand the set of cases for which it is known that all isomorphisms of Petit rings are monomial of type $G_{\tau,\alpha}$, when $\tau\in{\rm Aut}(S)$ commutes with $\sigma$. This was previously known in several cases, but so far only for isomorphisms, e.g. see [BP18, Theorems 28, 29]. The monomial isomorphisms $G_{\tau,\alpha,k}$ can be used to classify skew polycyclic codes, in particular skew constacyclic codes, which is why they are of particular interest (cf. Section 5.1).
The following results were originally proved only for $S_{0}$-isomorphisms for the case that $S$ is a field in [BP18, Theorems 28, 29], and generalized to the present setting in [Pum25]. To state them, set
@@
[Pum25, Theorem 4.1] Let $\alpha\in S^{\times}$ and assume that $\tau\in{\rm Aut}(S)$ commutes with $\sigma$. Then
$G_{\tau,\alpha}:\mathbb{S}_{f}\to\mathbb{S}_{h}$
is an isomorphism between nonassociative unital rings that restricts to the identity on $S_{0}^{\tau}$, if and only if
@@
When $S=K$ is a field, $\mathbb{K}_{f}$ and $\mathbb{K}_{h}$ are proper nonassociative Petit rings, and $\operatorname{Aut}(K)$ is abelian, we can say even more.
@@
[Pum25, Theorem 4.2] Let $\mathbf{m}\geq\mathbf{n}-1$ and assume that $f$ and $h$ do not generate two-sided ideals in $K[t;\sigma]$. Suppose $\operatorname{Aut}(K)$ is abelian. For any subfield $F\subset K_{0}$, the only ring isomorphisms
$G:\mathbb{K}_{f}\to\mathbb{K}_{h}$
that restrict to the identity on $F$ are of the form $G=G_{\tau,\alpha}$ such that $\tau\in\operatorname{Aut}(K)$, $K_{0}^{\tau}\subset F$, and $\alpha\in K^{\times}$ satisfying Theorem 2.2.
@@
2.4. Skew polycyclic and constacyclic codes
A linear code of length $\mathbf{n}$ over $S$ is a free submodule of the $S$-module $S^{\mathbf{n}}$. Let
$f(t)=t^{\mathbf{n}}-\sum_{i=0}^{\mathbf{n}-1}a_{i}t^{i}\in S[t;\sigma]$
be a reducible monic polynomial of degree $\mathbf{n}>1$. Then a linear code $C\subset S^{\mathbf{n}}$ is called a skew $(f,\sigma)$-polycyclic code, or skew polycyclic code for short, if for each $(c_{0},c_{1},\ldots,c_{\mathbf{n}-1})\in C$, also
$\left(0,\sigma(c_{0}),\sigma(c_{1}),\ldots,\sigma(c_{\mathbf{n}-2})\right)+\sigma(c_{\mathbf{n}-1})(a_{0},a_{1},\ldots,a_{\mathbf{n}-1})\in C.$
Note that
$\left(0,\sigma(c_{0}),\sigma(c_{1}),\ldots,\sigma(c_{\mathbf{n}-2})\right)+(\sigma(c_{\mathbf{n}-1})a_{0},\sigma(c_{\mathbf{n}-1})a_{1},\ldots,\sigma(c_{\mathbf{n}-1})a_{\mathbf{n}-1})$
$(\sigma(c_{\mathbf{n}-1})a_{0},\sigma(c_{0})+\sigma(c_{\mathbf{n}-1})a_{1},\sigma(c_{1})+\sigma(c_{\mathbf{n}-2})a_{2},\ldots,\sigma(c_{\mathbf{n}-2})+\sigma(c_{\mathbf{n}-1})a_{\mathbf{n}-1}).$
A skew $(f,\sigma)$-polycyclic code $C\subset S^{\mathbf{n}}$ is called a skew $(\sigma,a)$-constacyclic code, if $f(t)=t^{\mathbf{n}}-a$ for a non-zero $a\in S$, that is
$(c_{0},\dots,c_{\mathbf{n}-1})\in C\Rightarrow(\sigma(c_{\mathbf{n}-1})a,\sigma(c_{0}),\dots,\sigma(c_{\mathbf{n}-2}))\in C.$
As proven in [Pum25], the bijective map $\Phi:S^{\mathbf{n}}\rightarrow S[t;\sigma]_{\mathbf{n}}$,
$(c_{0},c_{1},\dots,c_{\mathbf{n}-1})\mapsto\sum_{i=0}^{\mathbf{n}-1}c_{i}t^{i},$
sends a skew $(f,\sigma)$-polycylic code $C$ of length $\mathbf{n}$ onto a principal left ideal $C(t)=\Phi(C)$ of the nonassociative Petit ring $\mathbb{S}_{f}=S[t;\sigma]/S[t;\sigma]f(t)$, the ambient ring of the code. That is, for a skew $(f,\sigma)$-polycylic code $C$ of length $\mathbf{n}$ over $S$ we denote by $C(t)$ the set of skew-polynomials $c(t)=\sum_{i=0}^{\mathbf{n}-1}c_{i}t^{i}$ associated to the codewords $(c_{0},\dots,c_{\mathbf{n}-1})\in C$, which is a left principal ideal in $\mathbb{S}_{f}$. Note that the left multiplication by $t$ with an element of a left principal ideal in the ambient ring $\mathbb{S}_{f}$ then produces exactly the shift that is used in the definition of a skew $(f,\sigma)$-polycyclic code.
It follows that any monomial homomorphism $G_{\alpha,\tau,k}$ of the ambient ring thus induces a Hamming-weight preserving isomorphism between codes.
3. Monomial homomorphisms from $S[t;\sigma]$ to a Petit ring
@@
Let $S$ be a commutative unital ring. We begin with a general restriction on monomial homomorphisms $G:S[t;\sigma]\to S[t;\sigma]$.
@@
Moreover, if $\tau$ and $\sigma$ commute, then $G$ is an isomorphism if and only if $k=1$ and $\alpha\in S^{\times}$.
@@
The cancellability of $\alpha$ implies that this holds for all $c\in S$ if and only if $\tau\sigma=\sigma^{k}\tau$. In particular, if $\sigma$ and $\tau$ commute, we must have $\sigma=\sigma^{k}$, or $k\equiv 1\mod\mathbf{m}$.
Suppose now that $\sigma$ and $\tau$ commute. If $G$ is to be an isomorphism, it must be surjective. For every nonzero $f\in S[t;\sigma]$, we have $\deg(G(t))=k\deg(f)$, hence all non-zero skew polynomials in the image of $G$ will have degree divisible by $k$, in particular so will $t$ which forces $k=1$. Moreover, if $\alpha$ were not invertible, then $t^{k}$ could not lie in the image of $G$. This implies that $\alpha\in S^{\times}$ and $k=1$.
Conversely, if $\sigma$ and $\tau$ commute then every $G$ with $\alpha\in S^{\times}$ and $k=1$ is clearly bijective. ∎
A special case of this result (where $\tau=\operatorname{id}$) has appeared as early as [Rim78, Theorem 3].
@@
Thus $G$ induces a ring isomorphism leaving $S_{0}^{\tau}$ fixed
$G:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b}\quad\text{where $\tau(a)=bN^{\sigma}_{\mathbf{n}}(\alpha)$.}$
In this section we show that these are all the Hamming-weight preserving homomorphisms from $S[t;\sigma]$ to $\mathbb{S}_{t^{\mathbf{n}}-b}$.
The first key obstruction for a map into $\mathbb{S}_{t^{\mathbf{n}}-b}=S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ to be a homomorphism is the power-associativity of the term $\alpha t^{k}$. We require a technical lemma that will in Theorem 3.3 be used to reduce the question of associativity of powers of $\alpha t^{k}$ in $\mathbb{S}_{t^{\mathbf{n}}-b}$ to a straightforward relation between $\alpha$ and $b$. Write $\lfloor x\rfloor$ for the greater integer that is less than or equal to $x$.
@@
We now demonstrate that the condition (3.1) is exactly the key to the power associativity of a monomial in $\mathbb{S}_{t^{\mathbf{n}}-b}$. Recall that an element $\mathsf{z}$ is said to be power-associative if $\mathsf{z}^{s}\mathsf{z}^{t}=\mathsf{z}^{t}\mathsf{z}^{s}$ for all $s,t\geq 1$, and this is equivalent to the statement that $\mathsf{z}^{s+t}$ is well-defined.
@@
in $\mathbb{S}_{t^{\mathbf{n}}-b}$. If $\mathsf{z}$ is power associative then
@@
Conversely, if (3.3) holds, then for every $s\in\mathbb{N}$, the expression $(\alpha t^{k})^{s}$ is well-defined in $\mathbb{S}_{t^{\mathbf{n}}-b}$ (that is, $\mathsf{z}$ is power associative) and we have
@@
We now prove the condition is necessary. Suppose $\mathsf{z}$ is power-associative. Then $\mathsf{z}\mathsf{z}^{r}=\mathsf{z}^{r}\mathsf{z}$, for $r>0$ the least integer such that $rk\geq\mathbf{n}$. We show that then (3.3) holds. We begin by noting that $\mathsf{z}^{r}$ is well-defined, since for all $s,\ell\geq 1$ such that $s+\ell=r$,
@@
In particular, Theorem 3.3 shows that the condition (3.3) is both necessary and sufficient for full power associativity of the element $\alpha t^{k}$ in $\mathbb{S}_{t^{\mathbf{n}}-b}$.
@@
$G=G_{\tau,\alpha,k}:S[t;\sigma]\to\mathbb{S}_{t^{\mathbf{n}}-b}$
given by $G(a)=\tau(a)$ and $G(t)=\alpha t^{k}$ induces a ring homomorphism leaving $S_{0}^{\tau}$ fixed if and only if
@@
Since $\tau$ commutes with $\sigma$ this holds if and only if $s-[sk]_{\mathbf{n}}\equiv 0\mod\mathbf{m}$; since $k\equiv 1\mod\mathbf{m}$ this implies $\mathbf{m}|\mathbf{n}$. Thus $\mathbf{n}=e\mathbf{m}$ for some positive integer $e$. Under this hypothesis, the condition $\alpha b=\sigma^{\mathbf{n}}(\alpha)\sigma^{k}(b)$ simplifies to $\alpha b=\sigma^{e\mathbf{m}}(\alpha)\sigma^{k}(b)=\alpha\sigma^{k}(b)$, since $\sigma$ has order $\mathbf{m}$ and implies that $\sigma(b)=b$ or $b\in S_{0}^{\times}$, and thus $\mathbb{S}_{t^{\mathbf{n}}-b}=S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is associative. Consequently the function $G=G_{\tau,\alpha,k}$ descends to a well-defined ring homomorphism $G:S[t;\sigma]\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$. ∎
@@
These conditions are quite restrictive, reflecting the constrained nature of these nonassociative algebras. For example, one deduces that even with $k=1$, the map $G_{\tau,\alpha,1}$ is a homomorphism from $S[t;\sigma]\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ only when $b\in S_{0}^{\times}$ and $\mathbf{m}|\mathbf{n}$. This is in fact to be expected: $S[t;\sigma]$ is associative whereas $\mathbb{S}_{t^{\mathbf{n}}-b}$ is associative if and only if these two strong conditions hold [Pet67, (15)].